# Exact Thresholds for Noisy Non-Adaptive Group Testing

### 10 Jan 2024 v1

@@ Abstract
In recent years, the mathematical limits and algorithmic bounds for probabilistic group testing having become increasingly well-understood, with exact asymptotic thresholds now being known in general scaling regimes for the noiseless setting. In the noisy setting where each test outcome is flipped with constant probability, there have been similar developments, but the overall understanding has lagged significantly behind the noiseless setting. In this paper, we substantially narrow this gap by deriving exact asymptotic thresholds for the noisy setting under two widely-studied random test designs: i.i.d. Bernoulli and near-constant tests-per-item. These thresholds are established by combining components of an existing information-theoretic threshold decoder with a novel analysis of maximum-likelihood decoding (upper bounds), and deriving a novel set of impossibility results by analyzing certain failure events for optimal maximum-likelihood decoding (lower bounds). Our results show that existing algorithmic upper bounds for the noisy setting are strictly suboptimal, and leave open the interesting question of whether our thresholds can be attained using computationally efficient algorithms.
 [^0] J. Chen is with the Department of Mathematics, The University of Hong Kong. The work was done when he was a visiting Ph.D. student in the Department of Computer Science, National University of Singapore (NUS). J. Scarlett is with the Department of Computer Science, Department of Mathematics, and Institute of Data Science, National University of Singapore (NUS). e-mail: chenjr58@connect.hku.hk; scarlett@comp.nus.edu.sg
@@ 1 Introduction
Group testing is a fundamental combinatorial and statistical problem with roots in medical testing and a variety of more recent applications including DNA testing, data storage, communication protocols, and COVID-19 testing. Alongside its uses in applications, group testing has had a long history of mathematical developments characterizing the number of tests required for reliable recovery, with increased interest in recent years [20, 6].
In particular, following a number of recent advances (which we describe in more detail in Section 1.2), a complete characterization is now available for the asymptotic number of tests required to achieve an error probability approaching zero in the noiseless setting [3, 39, 4, 29, 15, 17, 11]. Moreover, efficient algorithms are available whose performance matches these fundamental limits, with the Definite Defective (DD) algorithm [5, 29] or individual testing [2] sufficing in “less sparse” settings, and a spatial coupling approach [17] closing the remaining gaps in “more sparse” settings.
While similar advances have also been made in noisy group testing, its understanding has lagged significantly behind the noiseless setting. In particular, under the widely-considered symmetric noise model, optimal thresholds on the number of tests are only known in a narrow range of (sufficiently sparse) scaling regimes [39], and the degree of (sub-)optimality of practical algorithms is unknown [40, 41, 24].
In this paper, we address this gap by obtaining exact information-theoretic thresholds for the symmetric noise setting under two of the most commonly-considered test designs: i.i.d. Bernoulli and near-constant tests-per-item. Our results provide a significant improvement over the existing state-of-the-art, and demonstrate that all of the existing bounds for practical algorithms (to our knowledge) are suboptimal. We believe that this is a significant step forward in understanding noisy group testing, though we highlight that there are still two open problems that are beyond the scope of this work: A tight converse for general test designs, and an efficient algorithm that can attain the information-theoretic thresholds that we derive.
@@ 1.1 Problem Setup
$Y=\bigvee_{j\in S}X_{j}\oplus Z,$ (1)
[…]
@@
$P_{\mathrm{e}}:=\mathbb{P}[\widehat{S}\neq S].$ (2)
Here the probability is taken with respect to the uniformly random defective set, the possibly-randomized tests, and the noise. As was done in prior works such as [39, 17], we consider the goal of attaining $P_{\mathrm{e}}\to 0$ (with as few tests as possible), without seeking a precise convergence rate.
We will state our theorems in terms of bounds on $n$, but in our discussions and figures, we will instead use the notion of rate [6]:
@@
  • Under the Bernoulli design (e..g, see [33, 13, 3]), each item is independently placed in each test with probability $\frac{\nu}{k}$ for some $\nu>0$.
  • Under the near-constant weight design (e.g., see [29, 15]), each item is independently placed in $\Delta=\frac{\nu n}{k}$ tests chosen uniformly at random with replacement, for some $\nu>0$. (We leave integer rounding implicit, as it does not affect our results.) Since the same test may be chosen multiple times, the weight of each column in $\mathbf{X}$ may be (slightly) below $\Delta$.
@@ 1.2 Related Work
    Initial algorithmic bounds via the simple Combinatorial Orthogonal Matching Pursuit (COMP) algorithm with Bernoulli designs [13];
@@
    Improved algorithmic bounds via the more sophisticated (but still simple) Definite Defectives (DD) algorithm with Bernoulli designs [5];
@@
    A proof that the bounds in the preceding dot point are optimal among all designs, and a computationally efficient method for attaining this optimal threshold [17].
@@
Noisy group testing. In the noisy setting (with the symmetric noise model (1)), several counterparts to the above developments are available, but the list is significantly less complete:
@@
  • As an attempt towards #4, information-theoretic upper bounds were established in [41], but with tightness only shown for very small values of $\theta$.
@@
In this paper, we resolve the gaps in the entire list above for the noisy setting, except item #8.
In particular, we complete item #4 (precise-information-theoretic bounds; Bernoulli) which was only partially commenced in [39], and we complete items #5 (ensemble tightness; Bernoulli) and #7 (counterparts for the near-constant weight design), which have not been attempted before.
@@
  • For adaptive designs, near-optimal algorithms have long been known for the noiseless setting [27], and analogs were recently derived for the noisy setting [37, 43] and shown to significantly improve over the best non-adaptive results in denser (high $\theta$) regimes. We note that even following the improved non-adaptive bounds we derive in the present paper, significant improvements via adaptivity still remain.
  • Combinatorial non-adaptive group testing is a widely-studied problem [30, 22, 21, 20], and has also been studied under adversarial noise [14]. Under the condition of exact recovery, the stricter recovery criterion (namely, zero probability of error) increases the dependence on $k$ from linear to quadratic, meaning that it comes at a significant price in the number of tests. To our knowledge, precise constant factors have not been sought in this line of works.
  • If the exact recovery crtierion (2) is relaxed to approximate recovery, where $\alpha k$ false positives and false negatives are allowed in the reconstruction for some $\alpha>0$, then the study of noisy (and noiseless) probabilistic group testing becomes much simpler, with [39] giving upper and lower bounds that match in the limit of small $\alpha$. (With reference to the bound (4) below, the “difficult” second term disappears, and only the “simple” first term remains.) Analogous findings for the near-constant weight design can be inferred from [17, Sec. 5], and related studies of “all-or-nothing” phase transitions in group testing can be found in [44, 35].
@@ 1.3 Overview of Our Results
In this paper, we obtain exact information-theoretic thresholds for both the Bernoulli design and the near-constant weight design. To do so, we prove four results separately: Bernoulli achievability, Bernoulli converse, near-constant weight achievability, and near-constant-weight converse, where we use the term “converse” in a design-specific sense (i.e., ensemble tightness). We emphasize that all of our achievability results are based on a computationally intractable decoder that explicitly searches over all ($p\choose k$ possible defective sets; thus, we only prove achievability in an information-theoretic sense, and we defer efficient algorithms to future work.
@@
We note that both of the thresholds that we derive are somewhat complicated to state, but we will provide some intuition in Sections 2 and 3 regarding how the various terms arise. In all of our results, we make use of the design parameter $\nu>0$, noise level $\rho\in\big{(}0,\frac{1}{2}\big{)}$, and sparsity parameter $\theta\in(0,1)$, which is the constant such that $k=\Theta(p^{\theta})$. Since $\theta\in(0,1)$, we have $k\to\infty$ and $k=o(p)$.
Threshold for Bernoulli designs: The threshold for Bernoulli design with i.i.d. ${\rm Bernoulli}(\frac{\nu}{k})$ entries is given by
$n^{*}_{\rm Bern}=\max\left\{\frac{k\log\frac{p}{k}}{H_{2}(e^{-\nu}\star\rho)-H_{2}(\rho)},\frac{k\log\frac{p}{k}}{(1-\theta)\nu e^{-\nu}\min_{C>0,\zeta\in(0,1)}\max\{\frac{1}{\theta}f_{1}^{\rm Bern}(C,\zeta,\rho),f_{2}^{\rm Bern}(C,\zeta,\rho)\}}\right\},$ (4)
@@
$f_{1}^{\rm Bern}(C,\zeta,\rho)=C\log C-C+C\cdot D(\zeta\|\rho)+1,$ (5)
$f_{2}^{\rm Bern}(C,\zeta,\rho)=g^{\rm Bern}(C,\zeta,d^{*}_{\rm Bern},\rho),$ (6)
and where $g^{\rm Bern}(C,\zeta,\rho,d)$ and $d^{*}_{\rm Bern}$ are defined as
$g^{\rm Bern}(C,\zeta,d,\rho)=\rho d\log d+\big{(}\rho d-C(1-2\zeta)\big{)}\log\Big{(}\frac{\rho d-C(1-2\zeta)}{1-\rho}\Big{)}+1-2\rho d+C(1-2\zeta),$ (7)
$d^{*}_{\rm Bern}=\frac{C(1-2\zeta)+\sqrt{C^{2}(1-2\zeta)^{2}+4\rho(1-\rho)}}{2\rho},$ (8)
with the latter implicitly depending on $(C,\zeta,\rho)$. As we will argue in the proof, $d^{*}_{\rm Bern}$ is the minimizer of $g^{\rm Bern}$ with respect to $d$.
@@
Theorem 1.
(Achievability via Bernoulli Designs) Under the Bernoulli design with i.i.d. entries following ${\rm Bernoulli}(\frac{\nu}{k})$ for fixed $\nu>0$, there exists a decoding strategy such that $P_{e}\to 0$ as $p\to\infty$ provided that
$n\geq(1+\eta)n^{*}_{\rm Bern}$ (9)
@@
Theorem 2.
(Converse for Bernoulli Designs) Under the Bernoulli design with i.i.d. entries following ${\rm Bernoulli}(\frac{\nu}{k})$ for fixed $\nu>0$, any decoding strategy must have $P_{e}\to 1$ as $p\to\infty$ whenever
$n\leq(1-\eta)n^{*}_{\rm Bern}$ (10)
@@
Threshold for Near-Constant Weight Designs: The threshold for the near-constant weight design with $\Delta=\frac{\nu n}{k}$ placements per item is given by
$n^{*}_{\rm NC}=\max\left\{\frac{k\log\frac{p}{k}}{H_{2}(e^{-\nu}\star\rho)-H_{2}(\rho)},\frac{k\log\frac{p}{k}}{(1-\theta)\nu e^{-\nu}\min_{C\in(0,e^{\nu}),\zeta\in(0,1)}\max\{\frac{1}{\theta}f^{\rm NC}_{1}(C,\zeta,\rho,\nu),f^{\rm NC}_{2}(C,\zeta,\rho,\nu)\}}\right\},$ (11)
@@
$f_{1}^{\rm NC}(C,\zeta,\rho,\nu)=e^{\nu}D(Ce^{-\nu}\|e^{-\nu})+C\cdot D(\zeta\|\rho)$ (12)
$f_{2}^{\rm NC}(C,\zeta,\rho,\nu)=g^{\rm NC}(C,\zeta,d^{*}_{\rm NC},\rho,\nu),$ (13)
and where $g^{\rm NC}(C,\zeta,d,\rho,\nu)$ and $d^{*}_{\rm NC}$ are given by
$g^{\rm NC}(C,\zeta,d,\rho,\nu)=e^{\nu}\cdot D\big{(}de^{-\nu}\|e^{-\nu}\big{)}+d\cdot D\Big{(}\frac{1}{2}+\frac{C(1-2\zeta)}{2d}\big{\|}\rho\Big{)}$ (14)
$d^{*}_{\rm NC}=\operatorname*{arg\,min}_{d}\leavevmode\nobreak\ g(C,\zeta,d,\rho,\nu)\leavevmode\nobreak\ \leavevmode\nobreak\ \text{subject to }|C(1-2\zeta)|\leq d\leq e^{\nu}.$ (15)
Here $d^{*}_{\rm NC}$ implicitly depends on $(C,\zeta,\nu)$ and can in fact be expressed in closed form, but the expression is rather complicated so is omitted here (see (228) in Appendix E).
Theorem 3.
(Achievability via Near-Constant Weight Designs) Under the near-constant weight design where each item is placed $\Delta=\frac{\nu n}{k}$ tests uniformly at random with replacement with $\nu=\Theta(1)$, there exists a decoding strategy such that $P_{e}\to 0$ as $p\to\infty$ provided that
$n\geq(1+\eta)n^{*}_{\rm NC}$ (16)
@@
Theorem 4.
(Converse for Near-Constant Weight Designs) Under the near-constant weight design where each item is placed $\Delta=\frac{\nu n}{k}$ tests uniformly at random with replacement with $\nu=\Theta(1)$, any decoding strategy must have $P_{e}\to 1$ as $p\to\infty$ whenever $p\to\infty$ whenever
$n\leq(1-\eta)n^{*}_{\rm NC}$ (17)
@@
Discussion. We illustrate our results numerically using the notion of rate (measured in bits/test), as defined in (3). The bounds for both designs are illustrated in Figure 2, where we observe a substantial improvement over the best known existing bounds. The horizontal part at small $\theta$ in Figure 2 corresponds to the capacity of the binary symmetric channel (namely, $\log 2-H_{2}(\rho)$), and the corresponding impossibility result holds for arbitrary test designs and even adaptive algorithms (e.g., see [33, 10]). Thus, in this regime, we have particularly strong guarantees of asymptotic optimality. By comparison, for higher $\theta$, we have not established optimality with respect to arbitrary designs, but we have at least established ensemble tightness (and hence exact thresholds) for the two random designs under consideration.
Among the practical algorithms, the most promising one for being asymptotically optimal in some regimes (namely, $\theta$ close to one) is noisy DD, since DD is known to exhibit such optimality in the noiseless setting [3] and under the one-sided “Z” and “reverse Z” noise models [41, 24]. However, our results indicate that under symmetric noise, the existing noisy DD bounds are strictly subtoptimal for all $\theta\in(0,1)$ (at least for the $\rho$ values we considered). Hence, more sophisticated techniques appear to be needed to attain our thresholds in a computationally efficient manner, with the spatial coupling approach of [17] perhaps being a promising candidate.
@@
While the proofs of our main results all have highly technical components and are quite lenghty, the general approaches taken are conceptually simple. We provide conceptual outlines in Sections 2 and 3, and provide the technical details in the appendices, starting with a roadmap in Appendix A. The outlines will mostly be the same for the two designs, despite the differing technical details.

### 23 Oct 2024 v2

@@ Abstract
In recent years, the mathematical limits and algorithmic bounds for probabilistic group testing have become increasingly well-understood, with exact asymptotic thresholds now being known in general scaling regimes for the noiseless setting. In the noisy setting where each test outcome is flipped with constant probability, there have been similar developments, but the overall understanding has lagged significantly behind the noiseless setting. In this paper, we substantially narrow this gap by deriving exact asymptotic thresholds for the noisy setting under two widely-studied random test designs: i.i.d. Bernoulli and near-constant tests-per-item. These thresholds are established by combining components of an existing information-theoretic threshold decoder with a novel analysis of maximum-likelihood decoding (upper bounds), and deriving a novel set of impossibility results by analyzing certain failure events for optimal maximum-likelihood decoding (lower bounds).
 [^0] J. Chen is with the Department of Mathematics, The University of Hong Kong. The work was done when he was a visiting Ph.D. student in the Department of Computer Science, National University of Singapore (NUS). J. Scarlett is with the Department of Computer Science, Department of Mathematics, and Institute of Data Science, National University of Singapore (NUS). e-mail: chenjr58@connect.hku.hk; scarlett@comp.nus.edu.sg. This work was supported by the Singapore National Research Foundation (NRF) under grant number A-0008064-00-00.
@@ 1 Introduction
Group testing is a fundamental combinatorial and statistical problem with origins in medical testing and a variety of more recent applications including fault detection, DNA testing, data storage, communication protocols, and COVID-19 testing. The problem involves a collection of $n$ items, with a smaller subset of $k$ items being defective; here we adopt widely-used terminology from fault detection applications, but other terminology may be substituted in other applications (e.g., in medical testing, replace “items” by “individuals” and “defective” by “infected”). There is a mechanism by which items can be pooled together and tested in groups, and in the noiseless scenario, a test returns positive if it contains at least one defective item, and returns negative otherwise. The goal is to reliably identify the set of defective items using as few tests as possible.
Alongside its uses in applications, group testing has had a long history of mathematical developments characterizing the number of tests required for reliable recovery, with increased interest in recent years [20, 6]. In particular, following a number of recent advances (which we describe in more detail in Section 1.2), a complete characterization is now available for the asymptotic number of tests required to achieve an error probability approaching zero in the noiseless setting [3, 39, 4, 29, 15, 16, 11]. Moreover, efficient algorithms are available whose performance matches these fundamental limits, with the Definite Defective (DD) algorithm [5, 29] or individual testing [2] sufficing in “less sparse” settings, and a spatial coupling approach [16] closing the remaining gaps in “sparser” settings.
Similar advances have also been made in noisy group testing, in which some positive tests get flipped to negative and/or vice versa. However, the understanding of this setting has lagged significantly behind the noiseless setting. In particular, under the widely-considered symmetric noise model, optimal thresholds on the number of tests are only known in a narrow range of (sufficiently sparse) scaling regimes [39], and the degree of (sub-)optimality of practical algorithms is unknown [40, 41, 24].
In this paper, we address this gap by obtaining exact information-theoretic thresholds for the symmetric noise setting under two of the most commonly-considered test designs: i.i.d. Bernoulli and near-constant tests-per-item. Our results provide a significant improvement over the existing state-of-the-art, and demonstrate that all of the existing bounds for practical algorithms (to our knowledge) fall short of the information-theoretic thresholds. We believe that this is a significant step forward in understanding noisy group testing, though we highlight two important challenges that we do not attempt to address: Proving a tight converse for arbitrary test designs, and devising efficient algorithms whose required number of tests matches the information-theoretic thresholds.
@@ 1.1 Problem Setup
$Y=\Big{(}\bigvee_{j\in S}X_{j}\Big{)}\oplus Z,$ (1.1)
with noise term $Z\sim\mathrm{Bernoulli}(\rho)$ for some noise level $\rho\in\big{(}0,\frac{1}{2}\big{)}$, where $\oplus$ denotes modulo-2 addition. [^1] It would also be of interest to extend our analyses to asymmetric binary noise models (where $1\to 0$ and $0\to 1$ flips have different probabilities), but we stick to the symmetric noise due to its widespread consideration in previous works, and to avoid adding another layer of complexity in an already highly technical analysis. With $n$ tests, there are $n$ such test vectors $X^{(1)},\dotsc,X^{(n)}$, and we represent these via an $n\times p$ matrix $\mathbf{X}$. We only consider non-adaptive testing, in which the entire matrix $\mathbf{X}$ must be specified prior to observing any test results; such designs are often preferred due to allowing the implementation of the tests in parallel. The test outcomes $Y^{(1)},\dotsc,Y^{(n)}$ are represented via a length-$n$ vector $\mathbf{Y}$, and we will sometimes use $\mathbf{Z}$ to denote the corresponding vector of noise terms. We assume independent noise across tests, i.e., $Y^{(1)},\dotsc,Y^{(n)}$ are conditionally independent given $\mathbf{X}$.
@@
$P_{\mathrm{e}}:=\mathbb{P}[\widehat{S}\neq S].$ (1.2)
Here the probability is taken with respect to the uniformly random defective set, the possibly-randomized tests, and the noise. As was done in prior works such as [39, 16], we consider the goal of attaining $P_{\mathrm{e}}\to 0$ as $p\to\infty$ (with as few tests as possible). We will state our theorems in terms of bounds on the number of tests $n$, but in our discussions and figures, we will instead use the notion of rate [6]:
@@
  • Under the Bernoulli design (e.g, see [33, 13, 3]), each item is independently placed in each test with probability $\frac{\nu}{k}$ for some design parameter $\nu>0$ that dictates the density of 1s in $\mathbf{X}$.
  • Under the near-constant weight design (e.g., see [29, 15]), each item is independently placed in $\Delta=\frac{\nu n}{k}$ tests chosen uniformly at random with replacement, for some $\nu>0$. (We leave integer rounding implicit, as it does not affect our results.) Since the same test may be chosen multiple times, the weight of a given column in $\mathbf{X}$ may be (slightly) below $\Delta$.
@@ 1.2 Related Work
    Initial algorithmic bounds for Bernoulli designs via a simple algorithm that marks items in negative tests as non-defective an the rest as defective, sometimes referred to as the Combinatorial Orthogonal Matching Pursuit (COMP) algorithm [13];
@@
    Improved algorithmic bounds for Bernoulli designs via an algorithm that declares a given item as defective if and only if it appears in any test in which all other items are known to be non-defective, referred to as the Definite Defectives (DD) algorithm [5];
@@
    A proof that these bounds for the near-constant weight design are optimal among all designs, and a computationally efficient method for attaining this optimal threshold using a spatially coupled randomized design [16].
@@
Noisy group testing. In the noisy setting (with the symmetric noise model (1.1)), several counterparts to the above developments are available, but the list is significantly less complete:
@@
  • As an attempt towards #4, information-theoretic upper bounds were established in [41], but with tightness only shown for very small values of the sparsity parameter $\theta$.
@@
In this paper, we establish exact thresholds for the Bernoulli and near-constant weight designs
in the noisy setting, thus establishing counterparts to items #4, #5, and #7 in the above list.
@@
  • For adaptive designs, near-optimal algorithms have long been known for the noiseless setting [27], and analogs were recently derived for the noisy setting [37, 43] and shown to significantly improve over the best non-adaptive results in denser (high $\theta$) regimes. We note that even following the improved non-adaptive bounds that we derive in the present paper, significant improvements via adaptivity still remain.
  • Combinatorial non-adaptive group testing is a widely-studied problem [30, 22, 21, 20], and has also been studied under adversarial noise [14]. Under the condition of exact recovery, the stricter recovery criterion (with zero probability of error) increases the dependence on $k$ from linear to quadratic, meaning that it comes at a significant price in the number of tests. To our knowledge, precise constant factors have not been sought in this line of works.
  • If the exact recovery criterion (1.2) is relaxed to approximate recovery, where $\alpha k$ false positives and false negatives are allowed in the reconstruction for some $\alpha>0$, then the study of noisy (and noiseless) probabilistic group testing becomes much simpler, with [39] giving upper and lower bounds that match in the limit of small $\alpha$. (With reference to the bound (1.4) below, the “difficult” second term disappears, and only the “simple” first term remains.) Analogous findings for the near-constant weight design can be inferred from [16, Sec. 5], and related studies of “all-or-nothing” phase transitions in group testing can be found in [44, 35].
@@ 1.3 Overview of Our Results
In this paper, we obtain exact information-theoretic thresholds for both the Bernoulli design and the near-constant weight design. To do so, we prove four results separately: Bernoulli achievability, Bernoulli converse, near-constant weight achievability, and near-constant-weight converse, where we use the term “converse” in a design-specific sense (i.e., ensemble tightness). We emphasize that all of our achievability results are based on a computationally intractable decoder that explicitly searches over all ($p\choose k$ possible defective sets; thus, we only prove achievability in an information-theoretic sense, without requiring computational efficiency.
@@
We note that both of the thresholds that we derive are somewhat complicated to state, but we will provide some intuition in Sections 2 and 3 regarding how the various terms arise. In all of our results, we make use of the design parameter $\nu>0$, noise level $\rho\in\big{(}0,\frac{1}{2}\big{)}$, and sparsity parameter $\theta\in(0,1)$, which is the constant such that $k=\Theta(p^{\theta})$. (Since $\theta\in(0,1)$, we have $k\to\infty$ and $k=o(p)$.) We also give some brief hints on the meaning of other notation that will appear:
  • The thresholds both contain minimizations over two parameters denoted by $C$ and $\zeta$; we will see that $C$ is related to the number of tests a defective item is placed in without any other defectives, and $\zeta$ is related to how many of those tests are flipped.
  • The thresholds both contain quantities denoted by $f_{1}$ and $f_{2}$, which are related to the occurrence of “bad events” associated with defective items and non-defective items, respectively. Moreover, $f_{2}$ contains an additional optimization parameter $d$, which is related to how many “relevant” positive tests the “bad” non-defective item is placed in.
We again refer the reader to Sections 2 and 3 for more detailed intuition.
Threshold for Bernoulli designs: The threshold for Bernoulli design with i.i.d. ${\rm Bernoulli}(\frac{\nu}{k})$ entries is given as follows:
$n^{*}_{\rm Bern}=\max\left\{\frac{k\log\frac{p}{k}}{H_{2}(e^{-\nu}\star\rho)-H_{2}(\rho)},\frac{k\log\frac{p}{k}}{(1-\theta)\nu e^{-\nu}\min_{C>0,\zeta\in(0,1)}\max\{\frac{1}{\theta}f_{1}^{\rm Bern}(C,\zeta,\rho),f_{2}^{\rm Bern}(C,\zeta,\rho)\}}\right\},$ (1.4)
@@
$f_{1}^{\rm Bern}(C,\zeta,\rho)=C\log C-C+C\cdot D(\zeta\|\rho)+1,$ (1.5)
$f_{2}^{\rm Bern}(C,\zeta,\rho)=\min_{d\geq\max\{0,C(1-2\zeta)/\rho\}}g^{\rm Bern}(C,\zeta,d,\rho),$ (1.6)
and where $g^{\rm Bern}(C,\zeta,\rho,d)$ is defined as
$g^{\rm Bern}(C,\zeta,d,\rho)=\rho d\log d+\big{(}\rho d-C(1-2\zeta)\big{)}\log\Big{(}\frac{\rho d-C(1-2\zeta)}{1-\rho}\Big{)}+1-2\rho d+C(1-2\zeta).$ (1.7)
We will also show in the proof that the minimizing choice of $d$ in (1.6) has the following closed-form expression:
$d^{*}_{\rm Bern}(C,\zeta,\rho)=\frac{C(1-2\zeta)+\sqrt{C^{2}(1-2\zeta)^{2}+4\rho(1-\rho)}}{2\rho}.$ (1.8)
@@
Theorem 1.1.
(Achievability via Bernoulli Designs) Under the Bernoulli design with i.i.d. entries following ${\rm Bernoulli}(\frac{\nu}{k})$ for fixed $\nu>0$, there exists a decoding strategy such that $P_{e}\to 0$ as $p\to\infty$ with a number of tests satisfying
$n\leq(1+\eta)n^{*}_{\rm Bern}$ (1.9)
@@
Theorem 1.2.
(Converse for Bernoulli Designs) Under the Bernoulli design with i.i.d. entries following ${\rm Bernoulli}(\frac{\nu}{k})$ for fixed $\nu>0$, for any decoding strategy to have $P_{e}\not\to 1$ as $p\to\infty$ it must be the case that
$n\geq(1-\eta)n^{*}_{\rm Bern}$ (1.10)
@@
Threshold for Near-Constant Weight Designs: The threshold for the near-constant weight design with $\Delta=\frac{\nu n}{k}$ placements per item is given as follows:
$n^{*}_{\rm NC}=\max\left\{\frac{k\log\frac{p}{k}}{H_{2}(e^{-\nu}\star\rho)-H_{2}(\rho)},\frac{k\log\frac{p}{k}}{(1-\theta)\nu e^{-\nu}\min_{C\in(0,e^{\nu}),\zeta\in(0,1)}\max\{\frac{1}{\theta}f^{\rm NC}_{1}(C,\zeta,\rho,\nu),f^{\rm NC}_{2}(C,\zeta,\rho,\nu)\}}\right\},$ (1.11)
@@
$f_{1}^{\rm NC}(C,\zeta,\rho,\nu)=e^{\nu}D(Ce^{-\nu}\|e^{-\nu})+C\cdot D(\zeta\|\rho)$ (1.12)
$f_{2}^{\rm NC}(C,\zeta,\rho,\nu)=\min_{d\,:\,|C(1-2\zeta)|\leq d\leq e^{\nu}}g^{\rm NC}(C,\zeta,d,\rho,\nu),$ (1.13)
and where $g^{\rm NC}(C,\zeta,d,\rho,\nu)$ is given by
$g^{\rm NC}(C,\zeta,d,\rho,\nu)=e^{\nu}\cdot D\big{(}de^{-\nu}\|e^{-\nu}\big{)}+d\cdot D\Big{(}\frac{1}{2}+\frac{C(1-2\zeta)}{2d}\big{\|}\rho\Big{)}.$ (1.14)
We will also derive a closed-form expression for the value $d^{*}_{\rm NC}(C,\zeta,\nu)$ achieving the minimum in (1.13), but it is omitted here due to being rather complicated (see (E.53) in Appendix E).
Theorem 1.3.
(Achievability via Near-Constant Weight Designs) Under the near-constant weight design where each item is placed $\Delta=\frac{\nu n}{k}$ tests uniformly at random with replacement with $\nu=\Theta(1)$, there exists a decoding strategy such that $P_{e}\to 0$ as $p\to\infty$ with a number of tests satisfying
$n\leq(1+\eta)n^{*}_{\rm NC}$ (1.15)
@@
Theorem 1.4.
(Converse for Near-Constant Weight Designs) Under the near-constant weight design where each item is placed $\Delta=\frac{\nu n}{k}$ tests uniformly at random with replacement with $\nu=\Theta(1)$, for any decoding strategy to have $P_{e}\not\to 1$ as $p\to\infty$ it must be the case that
$n\geq(1-\eta)n^{*}_{\rm NC}$ (1.16)
@@
Discussion. We illustrate our results numerically using the notion of rate (measured in bits/test), as defined in (1.3). The bounds for both designs are illustrated in Figure 2, where we observe a substantial improvement over the best known existing bounds. The horizontal part at small $\theta$ in Figure 2 corresponds to the capacity of the binary symmetric channel (namely, $\log 2-H_{2}(\rho)$), and the corresponding impossibility result holds for arbitrary test designs and even adaptive algorithms (e.g., see [33, 10]). Thus, in this regime, we have particularly strong guarantees of asymptotic optimality. By comparison, for higher $\theta$, we have not established optimality with respect to arbitrary designs, but we have at least established ensemble tightness (and hence exact thresholds) for the two random designs under consideration.
Among the practical algorithms, the most promising one for being asymptotically optimal in some regimes (namely, $\theta$ close to one) is noisy DD, since DD is known to exhibit such optimality in the noiseless setting [3] and under the one-sided “Z” and “reverse Z” noise models [41, 24]. However, our results indicate that under symmetric noise, the existing noisy DD bounds fall short of the information-theoretic threshold for all $\theta\in(0,1)$ (at least for the $\rho$ values we considered).
@@
For the near-constant weight design, the unusual shape of the optimal $\nu$ curve in Figure 3 (Right) is explained as follows:
  • For small enough $\theta$, the first branch of the maximum in (1.11) is strictly dominant, and it is maximized by $\nu=\ln 2$, which corresponds to having half positive and half negative test outcomes.
  • For large enough $\theta$, the second branch of the maximum in (1.11) is strictly dominant, and the corresponding maximizer turns out to be a slowly decreasing function;
  • In between these two regimes, there is a narrow “transition region” in which $\nu$ is chosen to equate the two branches of the maximum.
A similar kind of transition region is observed for the Bernoulli design (analogous to the noiseless setting [4]), but in that case the maximizer remains at $\nu=1$ after the transition region.
@@
While the proofs of our main results all have highly technical components and are quite lengthy, the general approaches taken are conceptually simple. We provide conceptual outlines in Sections 2 and 3, and provide the technical details in the appendices, starting with a roadmap in Appendix A. The outlines will mostly be the same for the two designs, despite the differing technical details.
Note on concurrent work. The concurrent work of Coja-Oghlan et al. [18] established the same achievability threshold as that of Theorem 1.3 (stated in a different form) using a computationally efficient spatially coupled design, and also established Theorem 1.4. Thus, their work makes the major contribution of showing that $n^{*}_{\rm NC}$ is also attainable using a polynomial-time algorithm when spatial coupling is incorporated. Despite attaining the same threshold, the two sets of achievability analyses are very different, and we believe that there is significant interest in understanding the fundamental limits of each distinct test design; in this sense, their achievability result does not imply Theorem 1.3 nor vice versa, and the results for Bernoulli designs are unique to the present paper. We also note that asymmetric binary noise models are additionally handled in [18], whereas we focus on symmetric noise.